Determine algebraically whether the function is even, odd, or neither
F(x) = 3x^2 - 1

Answer by ikleyn(52788)

.
F(-x) = 3*(-x)^2 - 1 = 3*x^2 -1 = F(x),


hence, the function F(x) is even.